OpenStudy (anonymous):

find equation of circle with a center of (0,5) passing through the origin

OpenStudy (anonymous):

x^2 + (y+5)^2 = 25

OpenStudy (anonymous):

so it should be -5 not +5

OpenStudy (anonymous):

that's right... IMAO

OpenStudy (anonymous):

the circle passes through (0,0) the radius of the circle is the distance between the center (0,5) and (0,0) \[d=r=5\]\[\left( x-a \right)^{2}+\left( y-b \right)^{2}=r^{2}\]\[\left( x-0 \right)^{2}+\left( y-5 \right)^{2}=5^{2}\]\[x^{2}+\left( y-5 \right)^{2}=25\]